# DreamPBR: Text-driven Generation of High-resolution SVBRDF with Multi-modal Guidance

### v1

@@ 3.2. Physically-based material diffusion
The core component of the diffusion model is the denoising U-Net module (Ronneberger et al., 2015) which conditioned on timestep $t$. Following Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), our model employs a deterministic forward diffusion process $q(z_{t}|z_{t-1})$ to transform latent vectors $z$ towards an isotropic Gaussian distribution. The U-Net network is specifically trained to reverse the diffusion process $q(z_{t-1}|z_{t})$, iteratively denoising the Gaussian noise back into latent vectors . Adopting the strategy proposed by Rombach et al. (2022), we incorporates the text feature $\tau(T)\in\mathbb{R}^{M\times d_{\tau}}$ into the intermediate layer of UNet through a cross-attention mechanism $\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)\cdot V$, where $Q=W_{Q}^{i}\cdot\varphi_{i}\left(z_{t}\right),K=W_{K}^{i}\cdot\tau(T),V=W_{V}^{i}\tau(T)$, $\varphi_{i}\left(z_{t}\right)\in\mathbb{R}^{N\times d_{\epsilon}^{i}}$ represents an intermediate representation of the UNet $\epsilon_{\theta}$, and $W_{V}^{i}\in\mathbb{R}^{d\times d_{\epsilon}^{i}}$, $W_{Q}^{i}\in\mathbb{R}^{d\times d_{\tau}^{i}},W_{K}^{i}\in\mathbb{R}^{d\times d_{\tau}^{i}}$ are learnable projection matrices.
@@
    Continuity across boundaries. Unlike classic padding methods such as zero padding, which may introduce artificial edges, circular padding ensures boundary continuity. It wraps image content around both horizontal and vertical boundaries, providing seamless transition when tiling.
@@
    Pattern preservation. Circular padding mainly affects the boundary area of image, leaving the central area and overall texture patterns unchanged.
@@ 3.3. Render-aware SVBRDF decoder
For the rendering loss, we adopt the sampling scheme proposed by Deschaintre et al. (2018b) to render nine images per material map. The images include three images rendered with independently distant light and view directions, and six images using near-field mirrored view and lighting directions. The rendering loss yield desireable SVBRDF reconstructions, achieved by encouraging the training process to focus on minimizing errors in crucial material parameters rather than treating them with equal importance.
@@
To address this, we introduce a highlight-aware albedo decoder $\mathcal{D}_{P}$, which is finetuned on a synthetic shaded-to-albedo dataset, ensuring robust regularization to effectively minimize highlight artifacts in albedo maps. For each material sample in our SVBRDF dataset, we simulate various lighting conditions and viewpoints by randomly positioning point lights and cameras parallel to the material plane and then render SVBRDFs to reference shaded images by a physically-based renderer.
@@
High-resolution material maps are essential for achieving photorealistic renderings. However, due to the memory and performance constraints, current diffusion models typically generate images at a resolution of $512\times 512$, which falls short for high-quality production rendering.
@@
We fine-tune the Real-ESRGAN with material data, which is trained on purely synthetic data, to more effectively capture the high-frequency details of materials. We incorporate a rendering loss (similar to Equation 3) into the training of super-resolution module to ensure that the generated details contribute to high-frequency shading effects rather than visual artifacts. We should note that special care must be taken for normal maps during augmentation involving flipping and rotation. The directions stored in normal map must be adjusted consistently with the map orientation to ensure consistent knowledge about surface normals.
@@ 3.4. Multi-model control
We propose three control mudules for DreamPBR: Pixel Control, Style Control, and Shape Control. These modules are designed to be decoupled, allowing for flexible combinations of multiple controls.
@@ 3.4.1. Pixel Control
Spatial property guidance is widely used in material creation by artists. Our Pixel Control module $G_{P}$ takes spatial control maps $I_{P}$ as input, utilizing the ControlNet architecture (Zhang et al., 2023), to guide the generation of spatially-consistent SVBRDFs. It supports controlled generation under sketch guidance and allows for image-to-image material inpainting with binary mask.
Our material LDM, as described in Section 3.2, is adapted in material domain, enabling plausible material generation controlled by pre-trained ControlNet checkpoints, which are trained with 2D supervision.However, we found that fine-tuning pre-trained ControlNet with material data significantly improves both controllability and the quality of generated materials. Specifically, we initialize our ControlNet using the ControlNet 1.1 Scribble checkpoint and fine-tune it on our SVBRDFs dataset. For generate the sketch guidance, we employ Pidinet (Su et al., 2021) for extracting sketches from albedo maps.
@@ 3.4.2. Style Control
The Style Control module $G_{I}$ takes images prompt $I_{S}$ as input and extracts the style characteristics to guide material generation. Inspired by Ye et al. (2023), image prompts $I_{s}$ are first encoded into image features by CLIP’s image encoder, and then embedded to material LDM using a decoupled cross-attention adaptation module. Multimodal material generation can be achieved by accompanying image prompt with text prompt.
Style Control module is able to effectively capture the appearance properties and structural information from the input images, to generate realistic and coherent material maps. This functionality is particularly useful in scenarios where materials need to be created based on specific exemplar images, which is a frequent requirement in material design industry. The interaction of the Style Control module with the Shape Control module will be detailed in Section 3.4.3.
@@ 3.4.3. Shape Control
The Shape Control module $G_{S}$ takes a segmented 3D model $O_{s}=\{O,s\}$ ($s$ denotes the geometry segmentation) and an optional photo exemplar $I_{o}$ as input and automatically generates high-quality material maps for each segmentation. When provided with only a segmented 3D model and a basic text prompt, we leverage large language models(LLMs) such as ChatGPT (Achiam et al., 2023), to enrich the text descriptions for each segmentation. For instance, given a 3D chair model, the language model can generate diverse text descriptions for each part like seat, leg, and armrest, each featured varied design styles. Furthermore, integration with existing Pixel Control and Style Control modules supports enhanced SVBRDF generation, ensuring superior quality and detailed material characteristics.
Our model integrates the material transfer pipeline TMT (Hu et al., 2022d) to automatically assign diverse generated materials to 3D shape based on an image exemplar. The TMT pipeline involves two stages: firstly, translating color from exemplar image to the projection of 3D shape and vice versa for segmentation results; secondly, assigning materials to projected parts using a material classifier network, based on the translated image. Unlike Hu et al. (2022d), we do not rely on predefined material collections in material assignment. Instead, we use predicted material labels of TMT directly as text prompts and translated images as image prompts in the Style Control module, allowing high-quality SVBRDF generation for each part. The proposed algorithm offers two significant advantages over traditional material transfer models: it expands material diversity beyond limited predefined material collections and transfer not only color and category information but also comprehensive material attributes including styles and spatial structures from 2D exemplar to 3D shapes, leveraging the capabilities of our Style Control module.
@@ 4.1. Implementation Details
Figure 3. The generation results of DreamPBR under text-only conditions: We randomly sampled numerous materials with various types and wide tags, by the prompts, “a PBR material of [type], [tags]”. Not only can DreamPBR generates materials that match the descriptions, but also some out-of-domain materials are created as well such as brick of snow-covered bricks, plastic of childrens playground slide, and wall of street art graffiti.
@@
Figure 4. Pixel Control’s results with the same pattern but different materials. The binary images in first column are control conditoins of different sketches and the generated materials are on its right with certain patterns same as our given images, following its material properties such as the edge of bricks and the growth rings of wood.
@@
Figure 5. Pixel Control’s results with the same material but different patterns. The first column shows the decriptions of materials and the control patterns are in the lower right corner of the rendering images. Like Figure 4, the results have also shown greate consistency of materials and patterns.
@@
Figure 6. Style Control’s results with the same style but different materials. The stlyed images are given in the first column, and each description of material is below the image, which provide users with more artistic ways to design textures.
@@
Figure 7. Diverse sampling results under the same prompts. We evaluate the diversity with a same basic description (top) and a same detailed description (bottom) but different random seeds. Both of them show quite different patterns and textures although a same prompt is used.
@@
Figure 8. Splicing results of our tileable textures. The first 3 rows show the tileable resutls generated under only text descriptions and the last row shows the results with Pixel Control and Style Control, whose control conditions are attached to the edge of images. All the textures we generated in this figure show high tileabilities without artifacts.
@@
Figure 9. Inpainting reults. The original texture is shown in the upper left corner, which is generated by the prompt “a PBR material of fabric, floral cotton fabric”. By different tags (above each image) and regions to be inpainted (bright area in each image), we can manipulate user-specified areas in the texutres according to preferences such as changing a leaf to colorful flowers.
@@
Prompt Style Pixel Output SVBRDF
@@
Figure 10. Results of combined control. We combine descriptions of materials (the first column), styled image(the second column), and binary image(the third column) to control the generated textures. Under the descriptions of materials, the generated results have both the given pattern and style incorporated into them.
@@
Figure 11. Results of ShapeControl. we leverage a large language model(LLM) to describe the segemented chair legs and chair back, which is used to generate their textures with the help of DreamPBR. Besides text-only descriptions(left two figures), addtional user-specific controls such as RGB images (upper right) and binary images (lower right) are allowed for personalized design as well.
@@
Figure 13. Qualitative comparison on pixel guidance. We compare our PixelControl agianst Zhou et al. (2022) in leather (first 4 rows) and tile (last 4 rows). In some cases, TileGen fails to generate textures that match given patterns such as (column 1, row 4), (column 2, row 4), and (column 1, row 8). There are some artifacts in results to fit given patterns from Zhou et al. (2022) such as (column 1, row 2), (column 2, row 6), (column 3, row 6) and (column 3, row 8). In contrast, ours show better cosistency to pattern based on their natural materials.
@@
Figure 14. Qualitative and quantitative comparison on PBR decoder with rendering loss. We trained two PBR decoder with and without rendering loss. With rendering loss, the decoded textures show better consistency in both rendering images and SVBRDF (espciallly for normal map) visually and achieve the lowest LPIPS of rendering images and RMSE of SVBRDF maps as illustrated in the table.
@@
Figure 15. Qualitative and quantitative comparison on Super-Resolution module. We show local zoom rendering results with high-resolution textures($2048\times 2048$), low-resolution textures($512\times 512$), and textures from three super-resolution modules(Wang et al., 2021) with different training strategies: pre-trianed one, fine-tuned one without $\mathcal{L}_{\text{render}}$ and fine-tuned one with $\mathcal{L}_{\text{render}}$(ours). Our final methods produce details and textures more consistent with ground truth compared with incomplete methods as shown in five images at the bottom in the figure. With fine-tuning on our datasets and $\mathcal{L}_{\text{render}}$, the table below presents our final models give lowest LPIPS between rendering images with super-resolution textures and rendering image with real high-resolution textures, and lowest RMSE between super-resolution SVBRDF maps and ground truth.
@@
Figure 16. Qualitative and quantitative comparison on highlight-aware module (HA). As shown in 16(a), the original albedo decoder from VAE may generate data with highlight and highlight-aware module can de-highlight them to make better generation results. As shown in 16(b), For those inputs without highlight, there is tiny difference regardless of whether the highlight module is used or not. When inputing images with highlight, we can de-highlight them by our highligt-aware module. For images with and without highlight, the table below numerically reaches the same conclusion by comparing the L1, PSNR, and LPIPS between real albedo maps and de-highlighted albedo maps.
@@
Figure 17. Qualitative comparison on fine-tuning ControlNet. We evalute the pre-trianed ControlNet and our fine-tuned version based on the pre-trained one to employ pixel-guidance generation. The textures from pre-trained ControlNet (w/o ft) are more like natrual images rather than textures.
@@ 4.1.1. Dataset Generation
With respect to the paired data for training ControlNet, we utilized Pidinet  (Su et al., 2021) to extract sketches from the albedo maps as mentioned in Section 3.4.1.
@@ 4.1.2. Other Details
DreamPBR was trained on quadruple Nvidia RTX 3090 GPUs. During the training of Material LDM, we employ Adam as our optimizer with base learning rate of $1.6\times 10^{-3}$ and closed learning rate scaling. Starting with the stable-diffusion-v1-5 checkpoint for 9000 epochs, we finetune it approximately 10 days. For the training of the PBR decoder, we set the base learning rate to $4.5\times 10^{-6}$ and enabled scale_lr, taking 4 days totally in which the output channels of the decoder were set to 8, with albedo and normal having three channels each, and metallic and roughness being single-channel. For the highlight-aware albedo decoder, we set the base learning rate to $4.5\times 10^{-6}$ and enabled scale_lr, taking 2 days totally in which the output channels of the decoder were set to 3. We incorporate rendering loss as detailed in Section Section 3.3 during the training process above.

### v2

@@ 3.2. Physically-based material diffusion
The core component of diffusion model is the denoising U-Net module (Ronneberger et al., 2015) which is conditioned on timestep $t$. Following Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), our model employs a deterministic forward diffusion process $q(z_{t}|z_{t-1})$ to transform latent vectors $z$ towards an isotropic Gaussian distribution. The U-Net network is specifically trained to reverse the diffusion process $q(z_{t-1}|z_{t})$, iteratively denoising the Gaussian noise back into latent vectors. Adopting the strategy proposed by Rombach et al. (2022), we incorporates the text feature $\tau(T)\in\mathbb{R}^{M\times d_{\tau}}$ into the intermediate layer of UNet through a cross-attention mechanism $\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)\cdot V$, where $Q=W_{Q}^{i}\cdot\varphi_{i}\left(z_{t}\right),K=W_{K}^{i}\cdot\tau(T),V=W_{V}^{i}\tau(T)$, $\varphi_{i}\left(z_{t}\right)\in\mathbb{R}^{N\times d_{\epsilon}^{i}}$ represents an intermediate representation of the UNet $\epsilon_{\theta}$, and $W_{V}^{i}\in\mathbb{R}^{d\times d_{\epsilon}^{i}}$, $W_{Q}^{i}\in\mathbb{R}^{d\times d_{\tau}^{i}},W_{K}^{i}\in\mathbb{R}^{d\times d_{\tau}^{i}}$ are learnable projection matrices.
@@
    Continuity across boundaries. Unlike classic padding methods such as zero padding, which may introduce artificial edges, circular padding ensures boundary continuity. It wraps image content around both horizontal and vertical boundaries, providing a seamless transition when tiling.
@@
    Pattern preservation. Circular padding mainly affects the boundary area of the image, leaving the central area and overall texture patterns unchanged.
@@ 3.3. Render-aware SVBRDF decoder
For the rendering loss, we adopt the sampling scheme proposed by Deschaintre et al. (2018b) to render nine images per material map. The images include three images rendered with independently distant light and view directions, and six images using near-field mirrored view and lighting directions. The rendering loss yields desirable SVBRDF reconstructions, achieved by encouraging the training process to focus on minimizing errors in crucial material parameters rather than treating them with equal importance.
@@
To address this, we introduce a highlight-aware albedo decoder $\mathcal{D}_{P}$, which is finetuned on a synthetic shaded-to-albedo dataset, ensuring robust regularization to effectively minimize highlight artifacts in albedo maps. For each material sample in our SVBRDF dataset, we simulate various lighting conditions and viewpoints by randomly positioning point lights and cameras parallel to the material plane and then rendering SVBRDFs to reference shaded images by a physically-based renderer.
@@
High-resolution material maps are essential for achieving photorealistic renderings. However, due to the memory and performance constraints, current diffusion models typically generate images at a resolution of $512\times 512$, which falls short of high-quality production rendering.
@@
We fine-tune the Real-ESRGAN with material data, which is trained on purely synthetic data, to more effectively capture the high-frequency details of materials. We incorporate a rendering loss (similar to Equation 3) into the training of the super-resolution module to ensure that the generated details contribute to high-frequency shading effects rather than visual artifacts. We should note that special care must be taken for normal maps during augmentation involving flipping and rotation. The directions stored in a normal map must be adjusted consistently with the map orientation to ensure consistent knowledge about surface normals.
@@ 3.4. Multi-model control
We propose three control modules for DreamPBR: Pixel Control, Style Control, and Shape Control. These modules are designed to be decoupled, allowing for flexible combinations of multiple controls.
@@ 3.4.1. Pixel Control
Spatial property guidance is widely used in material creation by artists. Our Pixel Control module $G_{P}$ takes spatial control maps $I_{P}$ as input, utilizing the ControlNet architecture (Zhang et al., 2023), to guide the generation of spatially-consistent SVBRDFs. It supports controlled generation under sketch guidance and allows for image-to-image material inpainting with a binary mask.
Our material LDM, as described in Section 3.2, is adapted in the material domain, enabling plausible material generation controlled by pre-trained ControlNet checkpoints, which are trained with 2D supervision. However, we found that fine-tuning pre-trained ControlNet with material data significantly improves both the controllability and the quality of generated materials. Specifically, we initialize our ControlNet using the ControlNet 1.1 Scribble checkpoint and fine-tune it on our SVBRDFs dataset. To generate the sketch guidance, we employ Pidinet (Su et al., 2021) for extracting sketches from albedo maps.
@@ 3.4.2. Style Control
The Style Control module $G_{I}$ takes image prompt $I_{S}$ as input and extracts the style characteristics to guide material generation. Inspired by Ye et al. (2023), image prompts $I_{s}$ are first encoded into image features by CLIP’s image encoder, and then embedded into material LDM using a decoupled cross-attention adaptation module. Multimodal material generation can be achieved by accompanying the image prompt with a text prompt.
Style Control module can effectively capture the appearance properties and structural information from the input images, to generate realistic and coherent material maps. This functionality is particularly useful in scenarios where materials need to be created based on specific exemplar images, which is a frequent requirement in the material design industry. The interaction of the Style Control module with the Shape Control module will be detailed in Section 3.4.3.
@@ 3.4.3. Shape Control
The Shape Control module $G_{S}$ takes a segmented 3D model $O_{s}=\{O,s\}$ ($s$ denotes the geometry segmentation) and an optional photo exemplar $I_{o}$ as input and automatically generates high-quality material maps for each segmentation. When provided with only a segmented 3D model and a basic text prompt, we leverage large language models(LLMs) such as ChatGPT (Achiam et al., 2023), to enrich the text descriptions for each segmentation. For instance, given a 3D chair model, the language model can generate diverse text descriptions for each part like seat, leg, and armrest, each featuring varied design styles. Furthermore, integration with existing Pixel Control and Style Control modules supports enhanced SVBRDF generation, ensuring superior quality and detailed material characteristics.
Our model integrates the material transfer pipeline TMT (Hu et al., 2022d) to automatically assign diverse generated materials to 3D shapes based on an image exemplar. The TMT pipeline involves two stages: firstly, translating color from exemplar image to the projection of 3D shape and vice versa for segmentation results; secondly, assigning materials to projected parts using a material classifier network, based on the translated image. Unlike Hu et al. (2022d), we do not rely on predefined material collections in material assignment. Instead, we use predicted material labels of TMT directly as text prompts and translated images as image prompts in the Style Control module, allowing high-quality SVBRDF generation for each part. The proposed algorithm offers two significant advantages over traditional material transfer models: it expands material diversity beyond limited predefined material collections and transfers not only color and category information but also comprehensive material attributes including styles and spatial structures from 2D exemplar to 3D shapes, leveraging the capabilities of our Style Control module.
@@ 4.1. Implementation Details
Figure 3. The generation results of DreamPBR under text-only conditions: We randomly sampled numerous materials with various types and wide tags, by the prompts, “a PBR material of [type], [tags]”. Not only can DreamPBR generate materials that match the descriptions, but also some out-of-domain materials are created as well such as brick of snow-covered bricks, plastic of a children’s playground slide, and  wall of street art graffiti.
@@
Figure 4. Pixel Control’s results with the same pattern but different materials. The binary images in the first column are control conditions of different sketches and the generated materials are on their right with certain patterns same as our given images, following their material properties such as the edge of bricks and the growth rings of wood.
@@
Figure 5. Pixel Control’s results with the same material but different patterns. The first column shows the descriptions of materials and the control patterns are in the lower right corner of the rendering images. Like Figure 4, the results have also shown great consistency of materials and patterns.
@@
Figure 6. Style Control’s results with the same style but different materials. The styled images are given in the first column, and each description of the material is below the image, which provides users with more artistic ways to design textures.
@@
Figure 7. Diverse sampling results under the same prompts. We evaluate the diversity with the same basic description (top) and the same detailed description (bottom) but different random seeds. Both of them show quite different patterns and textures although the same prompt is used.
@@
Figure 8. Splicing results of our tileable textures. The first 3 rows show the tileable results generated under only text descriptions and the last row shows the results with Pixel Control and Style Control, whose control conditions are attached to the edge of images. All the textures we generated in this figure show high tile abilities without artifacts.
@@
Figure 9. Inpainting results. The original texture is shown in the upper left corner, which is generated by the prompt “a PBR material of the fabric, floral cotton fabric”. By different tags (above each image) and regions to be inpainted (bright areas in each image), we can manipulate user-specified areas in the textures according to preferences such as changing a leaf to colorful flowers.
@@
Prompt Style Pixel Render SVBRDF
@@
Figure 10. Results of combined control. We combine descriptions of materials (the first column), styled images (the second column), and binary images (the third column) to control the generated textures. Under the descriptions of materials, the generated results have both the given pattern and style incorporated into them.
@@
Figure 11. Results of ShapeControl. we leverage a large language model(LLM) to describe the segmented chair legs and chair back, which is used to generate their textures with the help of DreamPBR. Besides text-only descriptions(left two figures), additional user-specific controls such as RGB images (upper right) and binary images (lower right) are allowed for personalized design as well.
@@
Figure 13. Qualitative comparison on pixel guidance. We compare our PixelControl against Zhou et al. (2022) in leather (first 4 rows) and tile (last 4 rows). In some cases, TileGen fails to generate textures that match given patterns such as (column 1, row 4), (column 2, row 4), and (column 1, row 8). There are some artifacts in results to fit given patterns from Zhou et al. (2022) such as (column 1, row 2), (column 2, row 6), (column 3, row 6), and (column 3, row 8). In contrast, ours show better consistency to pattern based on their natural materials.
@@
Figure 14. Qualitative and quantitative comparison on PBR decoder with rendering loss. We trained two PBR decoders with and without rendering loss. With rendering loss, the decoded textures show better consistency in both rendering images and SVBRDF (especially for normal maps) visually and achieve the lowest LPIPS of rendering images and RMSE of SVBRDF maps as illustrated in the table.
@@
Figure 15. Qualitative and quantitative comparison on Super-Resolution module. We show local zoom rendering results with high-resolution textures($2048\times 2048$), low-resolution textures($512\times 512$), and textures from three super-resolution modules(Wang et al., 2021) with different training strategies: pre-trained one, fine-tuned one without $\mathcal{L}_{\text{render}}$ and fine-tuned one with $\mathcal{L}_{\text{render}}$(ours). Our final methods produce details and textures more consistent with ground truth compared with incomplete methods as shown in five images at the bottom of the figure. With fine-tuning our datasets and $\mathcal{L}_{\text{render}}$, the table below presents our final models giving the lowest LPIPS between rendering images with super-resolution textures and rendering images with real high-resolution textures, and lowest RMSE between super-resolution SVBRDF maps and ground truth.
@@
Figure 16. Qualitative and quantitative comparison on highlight-aware module (HA). As shown in 16(a), the original albedo decoder from VAE may generate data with highlight and the highlight-aware module can de-highlight them to make better generation results. As shown in 16(b), For those inputs without highlight, there is a tiny difference regardless of whether the highlight module is used or not. When inputting images with the highlight, we can de-highlight them with our highlight-aware module. For images with and without highlights, the table below numerically reaches the same conclusion by comparing the L1, PSNR, and LPIPS between real albedo maps and de-highlighted albedo maps.
@@
Figure 17. Qualitative comparison on fine-tuning ControlNet. We evaluate the pre-trained ControlNet and our fine-tuned version based on the pre-trained one to employ pixel-guidance generation. The textures from pre-trained ControlNet (w/o ft) are more like natural images rather than textures.
@@ 4.1.1. Dataset Generation
Concerning the paired data for training ControlNet, we utilized Pidinet  (Su et al., 2021) to extract sketches from the albedo maps as mentioned in Section 3.4.1.
@@ 4.1.2. Other Details
DreamPBR was trained on quadruple Nvidia RTX 3090 GPUs. During the training of Material LDM, we employ Adam as our optimizer with a base learning rate of $1.6\times 10^{-3}$ and closed learning rate scaling. Starting with the stable-diffusion-v1-5 checkpoint for 9000 epochs, we finetuned it for approximately 10 days. For the training of the PBR decoder, we set the base learning rate to $4.5\times 10^{-6}$ and enabled scale_lr, taking 4 days total in which the output channels of the decoder were set to 8, with albedo and normal having three channels each, and metallic and roughness being single-channel. For the highlight-aware albedo decoder, we set the base learning rate to $4.5\times 10^{-6}$ and enabled scale_lr, taking 2 days total in which the output channels of the decoder were set to 3. We incorporate rendering loss as detailed in Section Section 3.3 during the training process above.